If the concentration of Mg2+ ion in seawater is 1.21x103 mg/L, what OH- concentration is required to precipitate Mg(OH)2 ?

Ksp(Mg(OH)2)=5.6x10-12

-OH concentration must be greater than _____ M.
Answer #1

solution :- Concentrates af mg t= 1-21x10 mg/ = 1-219/2 = 1.21 La moya = 0.05042 mol [mg2+] = 5:042x102 moye [or] =? Ksp melo